# C3VDv2 - Colonoscopy 3D video dataset with enhanced realism

### v1

@@ Methods
where $c,d,$ and $e$ are elements of the stretch matrix and $(c_{x},c_{y})$ correspond to center shift. These intrinsic camera parameters are computed during a camera calibration measurement. An $8\times 11$ checkerboard with a square size of 10 mm was utilized as a calibration target. The MATLAB 2023b Camera Calibration toolbox (Mathworks) was used to compute the fisheye camera intrinsics, which resulted in a mean reprojection error of 0.43 pixels. The camera intrinsics are provided in camera_intrinsics.json file.
@@
The 2D-3D registration process aligns geometric features between the video frames and the 3D model. The optimization and loss function proposed in the original C3VD study were used. During this process, the virtual camera traversed the pre-defined camera trajectory, rendering depth maps from the 3D colon model at each keyframe pose. Within each registration iteration, the current transformation estimate (T) was applied to the 3D model. Subsequently, depth maps were rendered at camera poses corresponding to the extracted video keyframes. Edge maps were extracted from the depth images using the Canny edge detection algorithm. An edge similarity score $\mathcal{S}_{e}$ was used to quantify the similarity between the edge maps obtained from the video keyframes and the rendered depth maps. After blurring the edge maps with a Gaussian kernel, the edge similarity score between the rendered depth’s edge maps $E_{\text{Depth}}$ and the target edge maps extracted from video keyframes $E_{\text{RGB}}$ is computed as:
@@
where K is the number of keyframe pairs, H and W represent the image height and width, respectively, and the dot product is applied between corresponding pixels in the edge maps. This edge similarity term was incorporated into the overall objective function used to optimize T:
@@
The initial model transform estimate $\mathbf{T}_{\mathrm{init}}$ was obtained through a manual alignment process using the 3D model-video overlay tool provided in the C3VD code repository. The CMA-ES optimization then refined this initial estimate. A population size of 100 samples was used during the optimization process. The parameter search space for the transformation was bounded within ±0.1 radians for rotations and ±7.5 millimeters for translations relative to the initial transform. Additionally, the initial step size $\sigma$ was set to 0.1. In some cases, the search space and sigma were further reduced individually for finer alignment.
The 2D-3D registration algorithm leveraged the similarity of edge features between video keyframes and rendered depth maps derived from the 3D model. Edges in the depth maps predominantly correspond to anatomical landmarks, such as haustral folds, which were partially occluded by debris in the RGB frames. To mitigate this limitation, video sequences v2 and v3 were acquired from the same phantom before and after debris, using an identical robotic camera trajectory. Given the stationary phantom orientation between acquisitions and the robotic arm’s sub-millimeter repeatability error, the model transform calculated during v2 registration was applied directly to the debris-affected v3 sequence. This approach facilitated consistent model-video alignment despite partial visual obstruction.
To synchronize the clean (v2) and debris-filled (v3) colonoscopy videos, we computed the temporal offset that maximized the correlation between their motion trajectories. Specifically, we extracted the translation components from the camera pose logs of both videos and computed their correlation while sliding one trajectory over the other. The frame offset corresponding to the highest correlation was used to align the videos. Excess frames were subsequently trimmed, resulting in paired frames of clean and debris-filled colon from the two videos.
@@
The dataset comprises two distinct colon shapes (c1 and c2), each segmented into seven to eight anatomical regions, with each segment further characterized by four unique textures and colors (t1, t2, t3, and t4). Overall, eight full colon phantom models were created from sixty phantom segments.
For each phantom segment, up to four video sequences were recorded. The first two videos, v1 and v2, depict a clean colon environment with variations in camera trajectories and imaging settings. The third video, v3, incorporates fecal debris, blood, foam, and other colon artifacts to simulate realistic scenarios. Notably, the camera trajectory and imaging settings for v3 are identical to that of v2. These videos are stored in the Registered Videos folder. The fourth video for each segment (v4) features phantom deformation, with stationary or linear camera motion, and always includes debris. The v4 videos can be found in the Deformation folder. A total of 184 short videos were recorded, comprising 169 rigid (v1, v2 & v3) and 15 deformation (v4) videos with a total of 74,071 frames.
During video recording, parameters such as camera trajectory, speed, and edge enhancement settings were varied to simulate diverse colonoscopy scenarios. These variations, along with simulated artifacts and challenging cases, are comprehensively documented in the C3VDv2_Data_Summary_Sheet.xlsx, explained below.
@@
Following 2D-3D registration, ground truth maps were generated for each RGB video frame. The folder structure for each registered video is as follows:
@@
  • RGB Frame: rgb/NNNN.png represents the raw (distorted) video frame from the Olympus CF-HQ190L video colonoscope. The black border with video metadata was cropped, resulting in an image size of 1350 x 1080 pixels. NNNN denotes the 4-digit frame number within the video.
@@
  • Depth Frame: depth/NNNN_depth.tiff represents the depth along the camera frame’s Z-axis, clamped between 0 and 100 mm, and linearly scaled and encoded as a 16-bit grayscale image. For example, a pixel value of 16,384 corresponds to a depth of 25 mm.
@@
  • Surface Normal Frame: normals/NNNN_normals.tiff stores the X, Y, and Z components of the surface normal vector for each surfel in the R, G, and B color channels, respectively. Components are linearly scaled from $\pm$1 to 0-65535 and encoded as a 16-bit color image. Normal vector directions are defined with respect to the camera coordinate system: +x points right, +y points down, and +z points along the viewing direction (i.e., away from the camera).
@@
  • Optical Flow Frame: optical_flow/NNNN_flow.tiff depicts the optical flow from the current to the previous frame. X-direction motion (left to right, clamped between -20 to 20 pixels) is stored in the red channel, and Y-direction motion (up to down, clamped between -20 to 20 pixels) is stored in the green channel. Flow values are linearly scaled and encoded as a 16-bit color image.
@@
  • Occlusion Frame: occlusions/NNNN_occlusion.tiff indicates pixels that occlude other mesh faces within 100 mm of the camera origin, assigning a value of 255 to these pixels and 0 to all others. This binary data is encoded as an 8-bit grayscale image.
@@
  • Diffuse Frame: diffuse/NNNN_diffuse.png encodes Lambertian reflectance, computed using the dot product of the surface normal and the direction of the incident light. Reflectance values range from 0.1 to 1.0 and are linearly scaled and encoded as an 8-bit grayscale image.
@@
  • 3D Model and Coverage Map: coverage_mesh.obj stores the ground truth triangulated mesh. Texture vertices store coverage values, where vt=1 indicates an observed face, and vt=2 indicates an unobserved face.
We also include eight videos of simulated screening colonoscopy in the Simulated Screening Videos folder, comprising 95,300 frames. Each folder comprises all recorded RGB frames and a pose.txt file containing the camera poses in a frame-wise homogeneous format.
The C3VDv2 Data Summary Sheet contains the following columns:
@@
  • Quantitative Score: Comprises a combination of dice score (overlap) and chamfer distance (misalignment) on edges extracted from RGB frames and depth frames rendered from the 3D model. Refer to section Technical Validation for more details on this score.
Two calibration sequences are provided for geometric and photometric calibration in the Camera Calibration folder:
@@
  • camera_calib_vicalib.avi: This video features the “big_pattern” target from the Endomapper repository. The accompanying big_pattern.pdf file, when printed at 100% scale, includes a grid spacing of 5.29 mm, a large radius of 1.58 mm, and a small radius of 1.06 mm.
It is important to note that the pixel-wise depth, normals, occlusion maps, and 3D models for debris-filled colon videos (v3) are all determined from the pixel-wise ground truth for the ‘clean’ version of the colon without debris. These do not account for changes in depth, normals, or occlusion due to the addition of debris. The deformation videos (v4) and simulated screening videos do not include pixel-wise ground truth maps.
@@
C3VDv2 contains paired clean colon (v2) and debris-filled colon videos (v3), with identical camera trajectories and acquisition settings. To synchronize the videos, the time delay was calculated by maximizing the correlation between the translation components of the robotic arm poses. After trimming the beginnings and/or ends of the videos to account for this delay, the model transform from the clean colon video was applied to the debris-filled colon video. This reuse relies on two assumptions: the relative orientation of the model to the robotic arm remains constant between the two acquisitions, and the robotic arm has sub-millimeter pose repeatability. To validate these assumptions and the frame alignment, the similarity between paired frames was assessed using the Structural Similarity Index (SSIM). Ten random frames from the clean colon video were compared to a 31-frame window (offset +15 to -15) centered on the corresponding debris-filled colon frame. The average SSIM over the 10 frames was reported for each offset. The plot of average SSIM and standard deviation across all C3DVv2 video pairs (Figure 6a) shows a maximum at an offset of 0 frames (SSIM = 0.91), indicating maximum similarity between paired images. Importantly, the synchronization method only uses information from the camera poses, while the validation method is based only on comparing frames from the videos. This validates the synchronization method and justifies reusing the model transform for debris-filled videos.
2D-3D Registration Scoring
An edge-based 2D–3D registration method was utilized to align the 3D model with video frames by optimizing the model transformation. The registration algorithm was validated by Bobrow et al.[29] in C3VD using synthetic sequences. The proposed loss function demonstrated minimal rotational and translational error compared to alternative loss functions.
The 2D–3D registration performed well across most videos. However, we observed that the 3D model does not align optimally with the phantom in some cases, due to phantom manufacturing defects. Since the true transformation for real videos is unknown, a qualitative scoring system was implemented to assess the perceived alignment between the video and the 3D model. The qualitative alignment assessment was performed by manually reviewing video frames overlaid with their corresponding depth maps rendered from the 3D model.
@@
We note that pouring silicone into a mold sometimes results in bubbles partially embedded in the casted silicone surface. After curing, extra silicone was applied to patch these bubble holes; however, this can lead to an overflow outside the holes. Consequently, when such a phantom was placed in the mold for imaging, any excess silicone will cause local deformation of the phantom shape. This was the source for most of the videos with alignment scores of 3.
Additionally, quantitative scoring was performed by evaluating the alignment of edges extracted from the 3D model and the RGB frames from the videos. For each clean colon video, the edges from all RGB and depth frames—rendered from the 3D model—were obtained using the DexiNed model. These edges were binarized, and the Dice coefficient and the Chamfer distance between the edges extracted from depth frames and the edges extracted from RGB frames were computed and averaged across all video frames. Quantitative alignment scores $Q_{\text{align}}$ were computed by combining standardized z-scores of the Dice coefficient $D_{z}$ (edge similarity) and Chamfer distance $C_{z}$ (edge displacement). The final metric is defined as:
@@
A higher value of the quantitative alignment metric signifies better alignment. This formulation rewards high Dice scores (strong overlap) and penalizes high Chamfer distances (large misalignment). Both the Dice coefficient and the Chamfer distance are used, as the Dice coefficient is insensitive to the degree of misalignment in instances where edges do not overlap.
A comparative analysis of each video’s quantitative and qualitative metrics revealed a strong correlation, as illustrated in Figure 6b. We observed that most videos classified with a qualitative score of 1 exhibited higher quantitative alignment scores compared to those classified as 3. However, some Score 1 videos demonstrated low quantitative scores, often due to the presence of very few edges. In such cases, even minor misalignments within a subset of edges result in a notable decrease in the quantitative metric. Conversely, videos classified as Score 2 or 3, which exhibited high quantitative scores, generally maintained good alignment for most edges. However, they were categorized as Score 2 and 3 due to phantom deformations, which affected only a small subset of edges, thus allowing the quantitative score to remain high. Figure 6c shows the distribution of challenging scenarios across different scores.
@@
The dataset is available for download and preview on the project webpage. The dataset is available under the CC BY-NC-SA 4.0 license.
Discussion and Conclusion
C3VDv2 provides a comprehensive platform for the quantitative validation of 3D reconstruction algorithms such as SLAM[39, 40, 41] and its sub-components[42], and depth estimation[43, 44]. The dataset includes ground truth 3D models, camera poses, pixel-level depth, surface normals, optical flow, and occluded regions. These elements allow for precise evaluation of SLAM algorithms and their ability to identify unobserved regions during colonoscopy, with coverage maps serving as ground truth references. With the provided camera poses, C3VDv2 is also suitable for volumetric rendering based techniques such as neural radiance fields[8, 45] and Gaussian splatting[9, 46]. The wide variety of tissue texture and color in phantoms enables testing for domain generalization. Using Score 1 and 2 videos for tasks that require ground truth with high pixel-wise accuracy is recommended. In addition to camera pose estimation, Score 3 videos would be useful for training weakly supervised, semi- or self-supervised algorithms [47].
The dataset’s high-resolution 3D models and phantom-making protocols are open-sourced, enabling researchers to modify, 3D print colon molds, and cast their own phantoms for imaging. These protocols can be expanded to fabricate other tissue phantoms, such as other regions of the gastrointestinal tract like the esophagus and stomach, and to create organ structures for surgical datasets. C3VDv2 addresses scenarios where 3D reconstruction algorithms commonly fail, such as tracking loss on textureless colon surfaces, scope immersion in fluid or debris, and lens cleaning events. Additionally, the dataset includes trajectories designed to potentially debug reconstruction algorithms, including straight-line in-and-out motions, loops with identical starting and ending points, and mirrored trajectories with lens cleaning in the middle. Representative examples are shown in Figure 7. The simulated screening videos performed by an experienced gastroenterologist on complete colon phantoms offer realistic camera trajectories for validation of camera pose estimation methods. Short videos (v4) featuring simulated phantom deformations could be used for qualitative evaluation of deformable 3D reconstruction techniques.
The availability of debris-filled and corresponding clean colon videos with identical camera settings and trajectories allows researchers to assess the impact of debris on algorithm performance quantitatively. This also allows the dataset to be used for other clinically relevant tasks, such as artifact detection[48] and haustral fold detection [49] in the presence of debris. While the fecal debris has not been classified in terms of bowel prep, researchers could extend the dataset by labeling the Boston Bowel Preparatory Scale[50] to the videos to have a granular understanding of their algorithms’ performances with varying levels of debris. C3VDv2 videos and their paired ground truth frames could also be used to generate synthetic data [51, 52] for semi-supervised [53] or self-supervised learning. Since the dataset summary sheet describes the events occurring in the colonoscopy videos, the video-text pairs could aid training medical vision-language models[54]. C3VDv2 also holds potential for polyp detection and tracking. Some polyps have been strategically positioned to be occluded by haustral folds. Also, fecal debris could generate false-positive detections, presenting a challenging validation set for lesion detection algorithms.
Where pixel-level ground truth is available (v1, v2, and v3 videos), the models are constrained to be rigid and static. This contrasts with in vivo colonoscopy, where the colon is moving and deformable. Real colonic tissue has dense vasculature [55], and polyps have microscopic surface patterns, which are not replicated in the phantoms described here. These features may serve as key points for feature-tracking algorithms in real colonoscopy, but may fail in the case of phantoms.
Future datasets could include phantoms with tuned optical scattering and absorption properties to simulate realistic light-tissue interactions [56, 57, 58]. Phantoms with polyps suitable for simulated polypectomy [59] would enable the incorporation of therapeutic tools. Another future research area is generating pixel-wise ground truth and time-stamped 3D models for the deformable colon.
@@
The code used for the dataset is publicly available at https://github.com/DurrLab/C3VD

### v2

@@ Methods
where $c,d,$ and $e$ are elements of the stretch matrix and $(c_{x},c_{y})$ correspond to center shift. These intrinsic camera parameters were computed during a camera calibration measurement. An $8\times 11$ checkerboard with a square size of 10 mm was utilized as a calibration target. The MATLAB 2023b Camera Calibration toolbox (Mathworks) was used to compute the fisheye camera intrinsics, which resulted in a mean reprojection error of 0.43 pixels. The camera intrinsics are provided in camera_intrinsics.txt file.
@@
The 2D/3D registration aligns geometric features between the video frames and the 3D model. The optimization and loss function proposed in the original C3VD [19] study were used. During this process, the virtual camera traversed the pre-defined camera trajectory, rendering depth maps from the 3D colon model at each keyframe pose. At each registration iteration, the current transformation estimate ($\mathbf{T}$) was applied to the 3D model. Depth maps were then rendered at camera poses corresponding to the selected keyframes. Canny edge detection was used to extract edge maps from the rendered depth images. An edge similarity score $\mathcal{S}_{e}$ was used to quantify the similarity between the edge maps obtained from the video keyframes and the rendered depth maps. After blurring the edge maps with a Gaussian kernel, the edge similarity score between the rendered depth’s edge maps $E_{\text{Depth}}$ and the target edge maps extracted from video keyframes $E_{\text{RGB}}$ was computed as:
@@
where K is the number of keyframe pairs, H and W represent the image height and width, respectively, and the dot product is applied between corresponding pixels in the edge maps. The similarity score was incorporated into the optimization objective for estimating $\mathbf{T}$:
@@
The initial transform $\mathbf{T}_{\mathrm{init}}$ was manually estimated using the model-video overlay tool from the C3VD code repository. The CMA-ES optimization then refined this initial estimate. The optimization used a population size of 100. The parameter search space for the transformation was bounded within $\pm$0.1 radians for rotations and $\pm$7.5 millimeters for translations relative to the initial transform. Additionally, the initial step size $\sigma$ was set to 0.1. In some cases, the search space and sigma were further reduced for finer alignment.
The 2D/3D registration algorithm leveraged the similarity of edge features between video keyframes and rendered depth maps derived from the 3D model. Edges in the depth maps predominantly correspond to anatomical landmarks, such as haustral folds, which are often partially occluded by debris in the RGB frames. To mitigate this limitation, video sequences v2 and v3 were acquired from the same phantom before and after debris, using an identical robotic camera trajectory. Given the stationary phantom orientation between acquisitions and sub-millimeter repeatability of the robotic arm, the model transform calculated during v2 registration was reused for the debris-affected v3 sequence. This approach facilitated consistent model-video alignment despite partial visual obstruction.
To synchronize the clean (v2) and debris-filled (v3) colonoscopy videos, the temporal offset maximizing the correlation between their motion trajectories was calculated. Translation components were extracted from the camera pose logs of both videos, and their correlation was computed by sliding one trajectory over the other. The offset yielding the highest correlation was used to align the videos. Excess frames were then trimmed, resulting in paired frames of clean and debris-filled colon from the two videos.
@@
The dataset is hosted on the Johns Hopkins Research Data Repository at https://doi.org/10.7281/T1/JC64MK, and is also available for preview and download at the project webpage - https://durrlab.github.io/C3VDv2/. The dataset comprises two distinct colon shapes (c1 and c2), each segmented into seven to eight anatomical regions, with each segment further characterized by four unique textures and colors (t1, t2, t3, and t4). A total of eight full colon phantom models were created from sixty phantom segments.
C3VDv2 contains 192 videos with a total of 169,371 frames. It comprises three different types of video sequences:
  • Pixel-level GT Videos: registered_videos were acquired with a static, undeformed colon phantom and are provided with per-frame ground truth maps (depth, normals, optical flow, etc.). Up to three videos were recorded per phantom segment:
    • v1: clean colon with a baseline camera trajectory and imaging settings.
    • v2: clean colon with a different camera trajectory and imaging settings as v1.
    • v3: debris-filled colon using the same camera trajectory and imaging settings as v2.
    This category includes 169 short videos with a total of 67,886 frames.
  • Deformation Videos: deformation_videos consist of v4 videos featuring externally induced active phantom deformation, captured with either static or linear camera motion. All videos include debris. Each folder contains all recorded RGB frames and a corresponding pose.txt file (if camera is not stationary). The camera poses are in a frame-wise homogeneous format. This folder contains 15 short videos with a total of 6,185 frames.
  • Simulated Screening Videos: simulated_screening_videos comprise full-colon withdrawal sequences performed by a gastroenterologist to capture realistic camera motion. Similar to deformation videos, only RGB frames and camera poses in pose.txt are provided. A total of 8 videos are included, comprising 95,300 frames.
@@
Pixel-level ground truth was generated only for the rigid phantom videos (v1–v3) in the registered_videos folder. The folder structure for each registered video is as follows:
@@
  • RGB Frame: rgb/NNNN.png represents the raw (distorted, uncompressed) video frame from the Olympus CF-HQ190L video colonoscope. The black border with video metadata was cropped, resulting in an image size of 1350 x 1080 pixels. NNNN denotes the 4-digit frame number within the video.
@@
  • Depth Frame: depth/NNNN_depth.tiff represents the depth along the camera frame’s Z-axis, clamped between 0 and 100 mm, and linearly scaled and stored as a 16-bit grayscale image. For example, a pixel value of 16,384 corresponds to a depth of 25 mm.
@@
  • Surface Normal Frame: normals/NNNN_normals.tiff stores the X, Y, and Z components of the surface normal vector for each surfel in the R, G, and B color channels, respectively. Components are linearly scaled from $\pm$1 to 0-65535 and stored as a 16-bit color image. Normal vector directions are defined with respect to the camera coordinate system: +x points right, +y points down, and +z points into the camera.
@@
  • Optical Flow Frame: optical_flow/NNNN_flow.tiff depicts the optical flow from the current to the previous frame. X-direction motion (left to right, clamped between -20 to 20 pixels) is stored in the red channel, and Y-direction motion (up to down, clamped between -20 to 20 pixels) is stored in the green channel. Flow values are linearly scaled and stored as a 16-bit color image.
@@
  • Occlusion Frame: occlusions/NNNN_occlusion.tiff indicates pixels that occlude other mesh faces within 100 mm of the camera origin, assigning a value of 255 to these pixels and 0 to all others. This binary data is stored as an 8-bit grayscale image.
@@
  • Diffuse Frame: diffuse/NNNN_diffuse.png encodes Lambertian reflectance, computed using the dot product of the surface normal and the direction of the incident light. Reflectance values range from 0.1 to 1.0 and are linearly scaled and stored as an 8-bit grayscale image.
@@
  • 3D Model and Coverage Map: coverage_mesh.obj stores the ground truth triangulated mesh. Texture vertices (vt) store coverage values, where vt=1 indicates an observed face, and vt=2 indicates an unobserved face.
During video recording, parameters such as camera trajectory, speed, and edge enhancement settings were varied to simulate diverse colonoscopy scenarios. These variations, along with simulated artifacts and challenging cases, are comprehensively documented in the C3VDv2_Data_Summary_Sheet.xlsx. It contains the following columns:
@@
  • Quantitative Score: Comprises a combination of dice score (overlap) and chamfer distance (misalignment) on edges extracted from RGB frames and depth frames rendered from the 3D model. Refer to the technical validation section for more details on this score.
  • YouTube Preview URL ID: Provides the ID to append to https://www.youtube.com/watch?v= to access a preview of the video containing the RGB frames, depth, normal, optical flow maps, and camera trajectory.
  • Total Frames: Indicates the total number of frames in the video folder.
The spherical omnidirectional camera intrinsics are given in camera_intrinsics.txt. Additionally, two calibration sequences are provided for geometric and photometric calibration in the camera_calibration folder:
@@
  • camera_calib_vicalib.avi: This video features the “big_pattern” target from the Endomapper repository [15]. The big_pattern.pdf file, when printed at 100% scale, includes a grid spacing of 5.29 mm, a large radius of 1.58 mm, and a small radius of 1.06 mm.
@@
C3VDv2 contains paired clean colon (v2) and debris-filled colon videos (v3), with identical camera trajectories and acquisition settings. To synchronize the videos, the time delay was calculated by maximizing the correlation between the translation components of the robotic arm poses. After trimming the beginnings and/or ends of the videos to account for this delay, the model transform from the clean colon video was applied to the debris-filled colon video. This transform reuse relies on two assumptions: the relative orientation of the model to the robotic arm remained unchanged between the two acquisitions, and the robotic arm has sub-millimeter pose repeatability. To validate these assumptions and assess frame alignment, the similarity between paired frames was assessed using the Structural Similarity Index (SSIM). Ten randomly selected frames from the clean colon video were each compared against a 31-frame window (offsets from +15 to -15) centered on the corresponding debris-filled colon frame. The average SSIM over the 10 frames was reported for each offset. The plot of average SSIM and standard deviation across all C3DVv2 video pairs (figure 7a) shows a maximum at an offset of 0 frames (SSIM = 0.91), indicating maximum similarity between paired images. Notably, the synchronization method utilizes information only from the camera poses, whereas the validation method relies solely on comparing frames from the videos. This cross-modal validation using independent data sources confirms the accuracy of the synchronization method and justifies reusing the model transform for debris-filled videos.
2D/3D Registration Scoring
An edge-based 2D/3D registration method was utilized to align the 3D model with video frames by optimizing the model transformation. The registration algorithm was previously validated in C3VD [19] using synthetic sequences. The proposed loss function demonstrated lower rotational and translational error compared to alternative loss functions.
While the registration performed well across most videos, some showed suboptimal alignment between the model and the real phantom, primarily due to phantom manufacturing defects. Since the true transformation for real videos is unknown, a qualitative scoring system was implemented to assess the perceived alignment between the video and the 3D model. The qualitative alignment assessment was performed by manually reviewing video frames overlaid with their corresponding depth maps rendered from the 3D model.
@@
Most Score 3 cases stemmed from silicone bubbles during casting. We note that pouring silicone into a mold sometimes results in bubbles partially embedded in the casted silicone surface. After curing, extra silicone was applied to patch these bubble holes; however, this can lead to an overflow outside the holes. Consequently, when such a phantom was placed in the mold for imaging, any excess silicone would cause local deformation of the phantom shape.
Additionally, quantitative scoring was performed by evaluating the alignment of edges extracted from the 3D model and the RGB frames from the videos. For each clean colon video, edges were extracted from both rendered depth frames and raw RGB frames using the DexiNed model. These edges were binarized, and the Dice coefficient and the Chamfer distance between the edge maps were computed and averaged across all video frames. Quantitative alignment scores $Q_{\text{align}}$ were computed by combining standardized z-scores of the Dice coefficient $D_{z}$ (edge similarity) and Chamfer distance $C_{z}$ (edge displacement). The final metric is defined as:
@@
A higher value of the quantitative alignment metric signifies better alignment. This formulation rewards high Dice scores (strong overlap) and penalizes high Chamfer distances (large misalignment). The Dice coefficient is insensitive to the degree of misalignment when edges do not overlap; hence, both the Dice coefficient and the Chamfer distance are used.
A comparative analysis of each video’s quantitative and qualitative metrics revealed a strong correlation, as illustrated in figure 7b. We observed that most videos classified with a qualitative score of 1 exhibited higher quantitative alignment scores compared to those classified as 3. However, some Score 1 videos demonstrated low quantitative scores, often due to the presence of very few edges. In such cases, a minor misalignment of the few existing edges can disproportionately lower the score, even when the overall registration is qualitatively excellent. Conversely, videos classified as Score 2 or 3, which exhibited high quantitative scores, generally maintained good alignment for most edges. However, they were categorized as Score 2 and 3 due to phantom deformations, which affected only a small subset of edges, thus allowing the quantitative score to remain high. Figure 7c shows the distribution of challenging scenarios across different scores.
@@
It is important to note that all pixel-wise ground truth maps and 3D models for debris-filled colon videos (v3) are derived from the corresponding ‘clean’ version of the colon without debris. These do not account for changes in depth, normals, or occlusion due to the addition of debris. The deformation videos (v4) and simulated screening videos do not include pixel-wise ground truth. For tasks requiring high pixel-level accuracy, such as depth estimation or normal prediction, videos with scores 1 and 2 are recommended. Score 3 videos may be useful for training weakly supervised, semi-supervised, or self-supervised paradigms [58] in addition to pose estimation.
All pixel-wise ground truth was acquired from rigid, static colon models, which deviates from the deformable in vivo colon environment. Real colonic tissue has dense vasculature [59, 60], and polyps have microscopic surface patterns, which are not replicated in the phantoms used in the dataset. These features may serve as key points for feature-tracking algorithms in real colonoscopy, but may fail in the case of phantoms.
Data Availability
The dataset is available for download and preview on the project webpage https://durrlab.github.io/C3VDv2/. It is hosted publicly on the Johns Hopkins Research Data Repository at https://doi.org/10.7281/T1/JC64MK and released under the CC BY-NC-SA 4.0 license.
@@
The code used for the dataset is publicly available at https://github.com/DurrLab/C3VD.